OpenStudy (anonymous):

please help me to solve these problems. Write the indicated terms. Represent "half a given number" if the given number is represented by: a. 6 b. a c. 4a d. 4y+3 e. 2-3x

OpenStudy (phi):

did you try part a. ? what is 1/2 of 6 ?

OpenStudy (anonymous):

yeah my ans is 3,is that correct?

OpenStudy (phi):

yes. when you want 1/2 of 6 you write \[ \frac{1}{2} \cdot 6\] or \[ \frac{6}{2} \] then simplify here you get 3 if you want 1/2 of a, all you can do is write it down \[ \frac{a}{2} \] and that is as simple as it gets

OpenStudy (phi):

try c.

OpenStudy (anonymous):

heres my ans. 3, a\2, 2a, 2y+3, 1-3x..is that correct?

OpenStudy (phi):

the last 2 are not. you do the same rule. for 4y+3 \[ \frac{4y+3}{2} \] you can't really simplify it, but you can rewrite it. to see how this works, let's use numbers. if we had \[ \frac{4+2}{2} \] we know the top is 4+2 or just 6, and 1/2 of 6 is 3 we can rewrite the problem as \[ \frac{4}{2}+\frac{2}{2} \] if we simplify both fractions we get 2+1 and that gives us 3 so back to \[ \frac{4y+3}{2} \] we can rewrite it as \[ \frac{4y}{2} +\frac{3}{2}\] and you can simplify the first fraction, but not the second, to get \[ 2y +\frac{3}{2}\] I would not bother, as this is not any simpler than leaving it as \[ \frac{4y+3}{2} \]